OpenStudy (anonymous):

WILL FAN AND MEDAL Write an equation in slope intercept form Passes through (-8,7) with a slope of 1/4

OpenStudy (gabebae):

I can help

OpenStudy (gabebae):

Give me a minute :)

OpenStudy (anonymous):

Thanks

OpenStudy (gabebae):

If you were to graph it the line would pass through the y axis at (0, 9)

OpenStudy (hihi67):

@GabeBae how would he know that

OpenStudy (anonymous):

gotta point i flat out suck at math

OpenStudy (hihi67):

-8 = x, 7 = y correct?

OpenStudy (hihi67):

the function formula is y=mx+b

OpenStudy (hihi67):

you already have three of those values, the x, the y, and the m (which is the slope)

OpenStudy (hihi67):

if you rewrite it inserting your numbers it gives you -8 = 1/4 * 7 + b

OpenStudy (hihi67):

this is a very simple equation, and all you have to do is solve for b, as it's your only unknown variable

OpenStudy (hihi67):

do you know how to continue @GriffLovesPittsburgh

OpenStudy (gabebae):

Well, he named two points (-8, 7) take the slope go up 1 and right 4 until you hit the y-axis you'll hit it at (0, 9) To write it in slope intercept form y=mx+b m= the slope which is 1/4 and B= the y-intercept The answer would be y= 1/4x + 9

OpenStudy (anonymous):

thanks gabe

OpenStudy (gabebae):

Do you get it now?

OpenStudy (hihi67):

@GabeBae dat ish is wayy too complcated

OpenStudy (anonymous):

yea

OpenStudy (gabebae):

How is that complicated?

OpenStudy (hihi67):

i dont think he would know how to find the y intercept...

OpenStudy (anonymous):

i am now closing the question

OpenStudy (tkhunny):

Just to be complete... Write an equation in slope intercept form Passes through (-8,7) with a slope of 1/4 Point-Slope Form (since that is what we are given) (y-k) = m(x-h) We are given: h = -8 and k = 7 (y-7) = m(x+8) We are given: m = 1/4 (y-7) = (1/4)(x+8) That's it. That's the equation of the line. Find the Slope-Intercept form. This just means "solve for y" and clean up the right-hand side. Distributive property: (y-7) = (1/4)x+(1/4)8 = (1/4)x + 2 Add 7 y-7+7 = (1/4)x+(1/4)8 = (1/4)x + 2+7 y = (1/4)x + 9 And we are done. In the future, just keep tying and keep asking. You'll get it. NO giving up!!